Systems Staff Meeting Notes 03/04/02

Systems Staff Meeting Notes
March 4, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart
Announcements
  1. Bobbie’s mother is celebrating her 82nd birthday today. Rich’s mom will be 87 this week. It is also his 10th wedding anniversary.
  2. Geoffrey watched the Lone Star restaurant burn down.
  3. Cynthia thought “The X-Files” was good last night.
  4. Nagging problem with missing items – FP2K disks are gone. We can no longer leave items out on the worktable for pickup. Vault inventory – everything that comes out of there must be logged & tagged. Written security procedures & spot checks are necessary. Must increase user awareness – many do not keep track of software disks or documentation. We end up having to replace too many items. We may need to store items for pickup in the green lockers in room 10. Geoffrey will try to backup (burn) a copy of FP2K today.
  5. Take home test – Bill wants everyone in Systems to make a list of all of our servers – where they are located and what they do. List all of the operating systems shared over the server – there’s about 20. The main reason that Bill is asking us to do this is because in a few weeks there will be one day when Winston, Bobbie, Geoffrey, Bill, Will & Suzy will all be out for a day for security training on campus.
Review Minutes of 2/25/02
  1. Ann #1 – Bobbie’s grandson is still quiet.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Dells are still not allocated.
  2. SAN order still needs to go out – is being fine-tuned.
  3. More memory for the new Dells has been ordered.
  4. Epson/HP scanner problems – HP scanners do not work under W2K. Epson scanners – have to download a new driver – the ones they came with do not work with W2K.
Liaisons/Support Issues
  1. Kiosk mode – TN3270 – Geoffrey and Winston are working on it.
  2. Virus scanner updates – Trend has posted articles detailing how to do manual updates, if automatic updates fail, on their Web site. Can go through their site to download the updates. They do not offer any other kind of work-around to the problem.
  3. Steve Carrico – has had a machine in the vault for pick-up for the past 2 weeks. He says that he no longer has a liaison on staff to install it. Bill says he will address the situation.
Networking
  1. Power Outage – the power did not go off at any critical locations. Bill stayed late, the North tower finally lost power around 8:30 AM. The back circuits were still live. Bill wants the circuits mapped to prevent unnecessary shut-downs in the future.
  2. We lost a HD on NT2 during the power outage. Now we need to order 2 more hard drives – one for backup, one for the shelf.
  3. The proxy server did not come back up. It had clock problems.
  4. NT2R – department folders disappeared – about ½ Gig of information. The share was turned off. Will hypothesizes that we need to have NT2R set up to replicate. This was a new share. Failover & share – both servers need to know it was shared.
  5. Cluster admin – if an item is X-ed out, turn it back on. This happens during failover. Dependencies don’t initialize during failure. Even though data appeared to be gone, it is there – it hasn’t been lost.
  6. Public printing – PSNT did not re-establish the connection after the power outage. Rebooting will fix it. If power goes down, all public workstations need to reboot.
  7. Network browse function – Will is tinkering with it. NT2 backup domain controllers are set correctly. Masterbrowse was set to false, had to change it to true. It is still not seeing all of the machines.
  8. W2K server conversion – FrontPage extension is still holding us up.
  9. Central Defrag – on hold.
Software Projects Status
  1. DLC on hold (Ying).
  2. Modifications to the Trouble Ticket database have been done.
  3. Cynthia is working on an extension of a report for Personnel – OPS data. They want to expand the current application to view more details.
  4. Host Explorer – Hilton is working on it with Geoffrey.
  5. NT4 upgrade – it’s a backup domain controller. We have to reinstall it to get it to be an active directory member. Winston will do the reinstall on Wednesday.
  6. SSN purge – had to do a survey of Systems databases to determine which applications we currently have that store SSN. We’ll have to do some coding and migration to swap out the old SSN for new numbers issued via the upcoming Directory Services Project. The new system is slated to replace the Dean’s Menu, etc. Many copies of info currently exist on a person. They are trying to consolidate it all to one source. They will be supplying a new, 9-digit University ID. This system will effectively ban any department from storing SSN on an online system by the end of this calendar year.
  7. Cynthia demoed the ILL Content Site – David is OK with it. It is an operational application available over the Web. Now we have to get them to use the Kangaroo disk drives. They will scan and create documents in LAD, upload them on the Kangaroos, then bring them over here for encryption. Bill has logged a request for broadband access for LAD.
  8. DHP server for Serials – Hilton found an article on PHP language vulnerability. There is a patch available. Occurs in UNIX/LINUX environments. Unknown how it would effect NT.
DP Coordinator
  1. Students are out on Spring Break this week. Debra reminds all to listen up for the phones.
  2. Bill asks that we keep an eye out to see if we notice an increase of unavailable phone lines in Systems. If it seems to be a reoccurring problem, then we may need to add another phone line to the office.
  3. Geoffrey will be out Wednesday-Friday.


Back to the Systems Homepage

Last updated March 11, 2002
by Debra Harris
debharr@mail.uflib.ufl.edu